The scores of 40,000 county employees who sat for a recent civil service were normally distributed with μ = 70 and σ = 10.

The number of county employees who scored above a 90 is what?
Answer #1

Solution :

Given that ,

P(x > 90) = 1 - p( x< 90)

=1- p P[(x - \mu) / \sigma < (90 - 70) / 10]

=1- P(z < 2.00 )

= 1 - 0.9772

= 0.0228

= 40000 * 0.0228 = 912

The number of county employees = 912